[Brahms-l] Shift report 20040318 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Thu Mar 18 2004 - 08:43:09 EST
------------------------------------------------------------
Date: Thu Mar 18 08:43:08 2004, Shift: 08:00-16:00
Supervisor: PS, with: 
Log book#15 pages 90 - 91

Shift Summary:
------------------------------------------------------------
1. Beam since around 4a.m.. Dumped at 8:00 + ? a.m.
   Next shift is scheduled for the beam development.

2. run#     tr2      tr3
   11114     7k      24k
   11115     53k     126k
   11116     59k     171k
   11117     75k     268k
   run 11118 still continues untill the beam dumps.  

3. No detector problems.
   Pii4 and pii12 were hanging for a short while for several  
   times.
   
   
    
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Mar 18 08:43:22 2004

This archive was generated by hypermail 2.1.8 : Thu Mar 18 2004 - 08:43:43 EST